A cube hs a volume of 8 cubic meterer.if each side is doubled in length what will its new volume be in cubic meters?
Answer by geetha_rama(94)

Volume of cube = x^3 (where x is the length of the side of the cube)
8 m3 = x ^3
2^3 = x^3
=> x = 2 m
When side is doubled,
new length = 2*2 = 4 m
new volume = 4 ^ 3 = 64 m3
Hence new volume is 64 cubic meters